OpenStudy (anonymous):

4m-4=4m solve this equation. prefer an explanation.

OpenStudy (anonymous):

It's impossible, unless infinity is an answer.

OpenStudy (anonymous):

-4 \[\neq\] 0

OpenStudy (anonymous):

4m = 4m + 4 0 = 4

OpenStudy (anonymous):

Impossible

OpenStudy (amistre64):

we can start by getting rid of all the 4s; they are a scaled version of: m-1 = m

OpenStudy (amistre64):

then notice that nothing fits :)

OpenStudy (sasogeek):

4=0, where 4 is an unreal number

OpenStudy (amistre64):

of course this assumes that we are using standard mathematical notation

OpenStudy (anonymous):

So we have 4m - 4 = 4m. If we factor out a 4, we have: 4(m-1) = 4m. But this does not help. 4m-4 = 4m. Subtracting 4 m from both sides, we have: -4 = 0m Division by zero is not permitted, so we have 0 = -4 which without using the complex number system is impossible, so the answer for a typical algebra course would be either infinity or no solutions. Hope this helps. http://www.tutorsean.net